OpenStudy (anonymous):

i^15=-i? is this correct?

OpenStudy (anonymous):

I believe you devide by 4 and use the remainder?

OpenStudy (anonymous):

yes i think i^14 * i= -i

OpenStudy (anonymous):

@jayanta how come i^15 = i^12 + i^3?

OpenStudy (anonymous):

note: sorry its no plus(+) sign its multiplication(*)

OpenStudy (anonymous):

i see

OpenStudy (anonymous):